# UniMo: Universal Motion Correction For Medical Images without Network Retraining

### v1

@@ Abstract
In this paper, we introduce a Universal Motion Correction (UniMo) framework, leveraging deep neural networks to tackle the challenges of motion correction across diverse imaging modalities. Our approach employs advanced neural network architectures with equivariant filters, overcoming the limitations of current models that require iterative inference or retraining for new image modalities. UniMo enables one-time training on a single modality while maintaining high stability and adaptability for inference across multiple unseen image modalities. We developed a joint learning framework that integrates multimodal knowledge from both shape and images that faithfully improve motion correction accuracy despite image appearance variations. UniMo features a geometric deformation augmenter that enhances the robustness of global motion correction by addressing any local deformations whether they are caused by object deformations or geometric distortions, and also generates augmented data to improve the training process. Our experimental results, conducted on various datasets with four different image modalities, demonstrate that UniMo surpasses existing motion correction methods in terms of accuracy. By offering a comprehensive solution to motion correction, UniMo marks a significant advancement in medical imaging, especially in challenging applications with wide ranges of motion, such as fetal imaging. The code for this work is available online  [^1] https://github.com/IntelligentImaging/UNIMO/tree/main.
1 Introduction
@@ 2 Background
2.1 Rigid Motion Estimation via Equivariant Filters
@@
2.2 Deformation Correction via Deformable Registration
@@ 2.3 Joint Optimization
Such joint approaches have demonstrated enhanced accuracy and robustness for single image modalities [15, 14]. However, these frameworks fall short when image pairs exhibit significant intensity changes. The issue arises because both dissimilarity terms in Eq.(2.3) are highly dependent on the images themselves, increasing the likelihood that the optimization will be biased by intensity variations. This limitation motivated the development of a multimodal motion correction framework that integrates knowledge from both image intensities and shape to effectively reduce errors caused by image intensity variations.
3 Methodology
3.1 Multi-modal Motion Correction
@@
3.2 Multimodal Rigid Transformation for Motion Correction
@@
3.3 Network Design and Training
@@
In addition to estimating deformations using LDDMM [45], we provide a deep learning model that learns stationary velocity fields [51] while maintaining comparable model accuracy. Advanced predictive image registration models, including TransMorph [52] and DiffuseMorph [53], can be integrated into our framework.
@@ 4 Experimental Evaluation
5 Results & Discussion
5.1 Single Modality

### v2

@@ Abstract
In this paper, we introduce a Universal Motion Correction (UniMo) framework, leveraging deep neural networks to tackle the challenges of motion correction across diverse imaging modalities. Our approach employs advanced neural network architectures with equivariant filters, overcoming the limitations of current models that require iterative inference or retraining for new image modalities. UniMo enables one-time training on a single modality while maintaining high stability and adaptability for inference across multiple unseen image modalities. We developed a joint learning framework that integrates multimodal knowledge from both shape and images that faithfully improve motion correction accuracy despite image appearance variations. UniMo features a geometric deformation augmenter that enhances the robustness of global motion correction by addressing any local deformations whether they are caused by object deformations or geometric distortions, and also generates augmented data to improve the training process. Our experimental results, conducted on various datasets with four different image modalities, demonstrate that UniMo surpasses existing motion correction methods in terms of accuracy. By offering a comprehensive solution to motion correction, UniMo marks a significant advancement in medical imaging, especially in challenging applications with wide ranges of motion, such as fetal imaging. The code for this work is available online  [^1] https://github.com/IntelligentImaging/UNIMO.
Index Terms:
Deep Learning, Medical Imaging Analysis, Motion Compensation Analysis, Multi-modality Fusion, Rigid Transformation Estimation
I Introduction
@@ II Background
II-A Rigid Motion Estimation via Equivariant Filters
@@
II-B Deformation Correction via Deformable Registration
@@ II-C Joint Optimization
Such joint approaches have demonstrated enhanced accuracy and robustness for single image modalities [15, 14]. However, these frameworks fall short when image pairs exhibit significant intensity changes. The issue arises because both dissimilarity terms in Eq.(II-C) are highly dependent on the images themselves, increasing the likelihood that the optimization will be biased by intensity variations. This limitation motivated the development of a multimodal motion correction framework that integrates knowledge from both image intensities and shape to effectively reduce errors caused by image intensity variations.
III Methodology
III-A Multi-modal Motion Correction
@@
III-B Multimodal Rigid Transformation for Motion Correction
@@
III-C Network Design and Training
@@
In addition to estimating deformations using LDDMM [45], we provide a deep learning model that learns stationary velocity fields [51] while maintaining comparable model accuracy. Advanced predictive image registration models, including TransMorph [52] and DiffuseMorph [53], can be integrated into our framework. Analogous to our previous work [14], we extend UniMo to a spatio-temporal approach that incorporates time information to handle the motion among data sequences. Please check the reference [14] for more details.
@@ IV Experimental Evaluation
V Results & Discussion
V-A Single Modality